# Motion simulation of radio-labeled cells in whole-body positron emission tomography

### v1

@@ Abstract
Cell tracking is a subject of active research gathering great interest in medicine and biology. Positron emission tomography (PET) is well suited for tracking radio-labeled cells in vivo due to its exceptional sensitivity and whole-body capability. For validation, ground-truth data is desirable that realistically mimics the flow of cells in a clinical situation. This study develops a workflow (CeFloPS) for simulating moving radio-labeled cells in a human phantom. From the XCAT phantom, the blood vessels are reduced to nodal networks along which cells can move and distribute to organs and tissues. The movement is directed by the blood flow which is calculated in each node using the Hagen-Poiseuille equation and Kirchhoff’s laws assuming laminar flow. Organs are voxelized and movement of cells from artery entry to vein exit is generated via a biased 3D random walk. The probabilities of whether cells move or stay in tissues are derived from rate constants of physiologically based compartment modeling. PET listmode data is generated using the Monte-Carlo simulation framework GATE based on the definition of a large-body PET scanner with cell paths as moving radioactive sources and the XCAT phantom providing attenuation data. From the flow simulation of 10000 cells, 100 sample cells were further processed by GATE and listmode data was reconstructed into images for comparison. As demonstrated by comparisons of simulated and reconstructed cell distributions, CeFloPS can realistically simulate the cell behavior of whole-body PET providing valuable data for development and validation of cell tracking algorithms.
@@ 1 Introduction
Cell imaging and tracking is a subject of active research gathering great interest in the medical and biomedical science community. Non-invasive imaging techniques using fluorescent or luminescent markers in optical imaging (OI), iron oxide particles in magnetic resonance imaging (MRI), or radio-labelled tracers in single photon emission computed tomography (SPECT) and positron emission tomography (PET) have been frequently used in preclinical studies to permit in vivo visualization of living cells [1]. For human application, MRI, PET and SPECT are the most promising tools as they provide either high spatial resolution (MRI) or excellent whole-body sensitivity (PET, SPECT). Of these imaging modalities, PET has become of special interest following labelled immune cells in the context of stem cell imaging and cell-based therapies [2].
@@
PET is particularly advantageous because it combines exceptional sensitivity with the possibility of whole-body coverage which has been further advanced in recent years by the design of large-body PET systems with extended coverage of the axial field-of-view in the range of 1 to 2 meters, improved time-of-flight capabilities and excellent system sensitivities [3]. Using either long-lived isotopes or reporter genes, immunoPET has great potential to identify biomarkers allowing to treat patients based on their most likely response to a dedicated immunotherapy [4, 5, 6].
Following the idea of PET-based particle tracking developed originally for technical applications, the outstanding sensitivity of PET was recently used to track radio-labeled singular cells [7, 8, 9, 10]. In this context, a new image reconstruction algorithm for PET was introduced by our group using an optimal transport based regularization technique [11] showing great potential for tracking single and multiple radio-labeled cells within a human body scenario.
@@
The current study is concerned with the development of a software workflow to accurately simulate the movement of radio-labeled cells within the human body in a clinical setting. This workflow generates dynamic cell flow data, which is then used in Monte Carlo simulations to model the underlying PET physics. Since the location and motion of the cells are known at all times, the cell flow PET simulation (CeFloPS) serves as ground-truth data for the development and validation of novel image reconstruction and tracking tools.
@@ 2.1 Simulation basis
CeFloPS is based on a special version of the XCAT phantom with high resolution 3D models of all organs, blood vessels, bones and muscles (Fig. 1a) in the form of STEP files [12, 13]. Using Inventor (Version 2023, Autodesk Inc., San Rafael, USA), these STEP files were converted into StL files (StereoLithography file format), providing triangulated surface meshes of all organs and structures. The StL files were further processed in Python (Version 3.11.5, Python Software Foundation) with the help of the trimesh library [14], allowing for operations on the meshes and giving access to the coordinates of the vertices. For further processing, meshes of all muscles and internal organs were converted into voxelized volumes with voxel size 0.5 mm $\times$ 0.5 mm $\times$ 0.5 mm, while the complex anatomy of blood vessel meshes was simplified by reducing it to node networks. The nodes describe the possible paths a cell can take and are placed along the central axis of a blood vessel, storing information about (i) its spatial coordinates, (ii) which nodes are adjacent in the direction of blood flow and (iii) the radius of the blood vessel.
@@
Figure 1: Mesh representation of the XCAT phantom defining vessel structures and organs (a). Only the larger vessels are defined by XCAT showing partly missing connections (b) and a just coarsely defined capillary bed (c).
@@ 2.2 Blood flow simulation
At a link, the probability $p_{\mathrm{node}}$ with which a cell will move to the next possible node is calculated based on the blood flow $\overline{Q}_{\mathrm{link}}$ in the link and $\overline{Q}_{\mathrm{node},i}$ in each following node:
$p_{\mathrm{node},i}=\frac{\overline{Q}_{\mathrm{node},i}}{\overline{Q}_{\mathrm{link}}}\quad\mathrm{with}\quad\overline{Q}_{\mathrm{link}}=\sum_{i=1}^{n}\overline{Q}_{\mathrm{node},i}$ (5)
@@
$\frac{\mathrm{d}Q(x)}{\mathrm{d}x}=2\pi xv(x)\,.$ (7)
Division by the mean blood flow rate yields the probability density function (PDF)
[…]
@@
$F_{p}(x)=\int_{0}^{x}f_{p}(t)\mathrm{d}t=\frac{x^{2}}{r^{2}}\left(2-\frac{x^{2}}{r^{2}}\right)\,.$ (9)
@@ 2.3 Organ distribution simulation
The random walk is initiated at a start-voxel, where a cell can move to one of the surrounding voxels. This gives 26 possible directions (provided the start-voxel does not reside on the organ’s periphery), all with the same probability. For the proceeding steps, the direction is biased in accordance with the previously determined weightings of the direction vectors in each voxel. If a cell reaches an end-voxel, the random walk concludes, followed by the cell exiting via the vein back into the nodal network for further blood flow simulation.
@@
Besides the biased random walk, a compartment model was implemented in the organ distribution simulation which governs the duration a cell stays in an organ. In each organ voxel a cell can be located in different virtual compartments. The exchange between compartments is determined by rate constants $k_{i}$ which also define the probabilities of whether a cell moves during the next time step or not. The number and structure of these compartments as well as the values of the rate constants are depending on the pharmacokinetics and physiology of the cell and the organ. This information must be known in advance to the CeFloPS simulation.
@@
Figure 4: The movement of cells within each organ is regulated by compartmental models, which determine the probability of a cell undergoing a random walk or being constrained to a virtual compartment.
As a first example and as proof-of-principle, we implemented a two-compartment 18F-Fluordesoxyglucose (FDG) model as shown in Fig. 4. The simulated cells in this example therefore do not behave like real, existing cells, but like an FDG tracer. For this two-compartment model, $C_{1}$ and $C_{2}$ denote the unbound and bound states of an FDG molecule in tissue and $C_{A}$ the free FDG molecule in the blood supply of the tissue. As long as cells are located in $C_{A}$ they undergo the random walk. If they enter $C_{1}$ or $C_{2}$, they stay at their position and do not move. The change in amount of cells in these compartments in dependency on the (arterial) blood $C_{A}$ can be described by the differential equations
$\begin{split}\biggl{[}\frac{\mathrm{d}}{\mathrm{d}t}C_{A}(t)&=-K_{1}C_{A}(t)+k_{2}C_{1}(t)\biggr{]}\,,\\ \frac{\mathrm{d}}{\mathrm{d}t}C_{1}(t)&=K_{1}C_{A}(t)-(k_{2}+k_{3})C_{1}(t)\quad\mathrm{and}\\ \frac{\mathrm{d}}{\mathrm{d}t}C_{2}(t)&=k_{3}C_{1}(t)\,.\end{split}$ (10)
The differential equation for $C_{A}$ is put into brackets as usually $C_{A}$ is a measured quantity and not a compartment like $C_{1}$ and $C_{2}$. For a better understanding of how the movement probabilities of a cell are derived, we still list the equations for $C_{A}$ as a compartment. In this example, the values of the rate constants $K_{1}$, $k_{2}$, and $k_{3}$ are taken from [18] and listed in Table 1.
@@
Table 1: Values of rate constants taken from [18] used in the simulation
@@
lung 0.023 0.205 0.001
liver 0.660 0.765 0.002
spleen 1.593 2.867 0.006
gray matter 0.107 0.165 0.067
myocardium 0.832 2.651 0.099
muscle 0.026 0.249 0.016
resta 0.553 1.213 0.039
a For organs not included in [18] the rate constants were set to the values of ”all regions”.
The differential equations in (10) can be analytically solved, yielding the following solutions for $C_{A}$ and $C_{1}$:
$\begin{split}\biggl{[}C_{A}(t)&=k_{2}C_{1}(t)*\mathrm{e}^{-K_{1}t}\biggr{]}\,,\\ C_{1}(t)&=K_{1}C_{A}(t)*\mathrm{e}^{-(k_{2}+k_{3})t}\,,\end{split}$ (11)
where * denotes the convolution operator. Since the solutions in (11) are linear combinations of exponential functions, we describe the time-dependent probability of transitions between compartments by the following equations based on the CDFs of exponential distributions:
$\begin{split}p_{C_{A}\rightarrow C_{1}}(t)&=1-\mathrm{e}^{-K_{1}t}\,,\\ p_{C_{1}\rightarrow C_{A}}(t)&=\frac{k_{2}}{k_{2}+k_{3}}\left(1-\mathrm{e}^{-\left(k_{2}+k_{3}\right)t}\right)\,\mathrm{and}\\ p_{C_{1}\rightarrow C_{2}}(t)&=\frac{k_{3}}{k_{2}+k_{3}}\left(1-\mathrm{e}^{-\left(k_{2}+k_{3}\right)t}\right)\,.\end{split}$ (12)
The equations in (12) indicate the probability of a transition between two compartments occurring from the beginning of the simulation up to time t. Based on these equations, a random decision is made at each time step as to whether a cell transitions or not. Cells entering a tissue are first located in the blood compartment $C_{A}$. As long as cells are in $C_{A}$ they undergo the (biased) random walk in the organ. Cells that have transitioned into $C_{1}$ do not partake in the random walk but stay at their position. For these cells, the decision is made after each step as to whether they go back to $C_{A}$ and are able to move again or whether they stay in $C_{1}$ or transition to $C_{2}$. A cell that transitions from $C_{1}$ back to $C_{A}$ is considered to be in the venous blood and undergoes random walk until reaching an end-voxel, without the possibility to reenter $C_{1}$. If cells enter $C_{2}$, they remain trapped in their position for the rest of the simulation and do no longer move.
For the FDG simulation, the paths and distributions of 10000 cells are simulated in parallel. For every time step, the organ and compartment in which each cell is located is recorded. This allows the generation of time profiles of cells in tissues. For a given tissue, the tissue concentration of cells per milliliter or the corresponding cell number (as the sum in $C_{1}$ and $C_{2}$) is calculated. The number of cells in the blood is made up of all cells that are still in the node network or in a $C_{A}$ compartment of an organ.

### v2

@@ Abstract
Cell tracking is a subject of active research gathering great interest in medicine and biology. Positron emission tomography (PET) is well suited for tracking radio-labeled cells in vivo due to its exceptional sensitivity and whole-body capability. For validation, ground-truth data is desirable that realistically mimics the flow of cells in a clinical situation. This study develops a workflow (CeFloPS) for simulating moving radio-labeled cells in a human phantom. From the XCAT phantom, the blood vessels are reduced to nodal networks along which cells can move and distribute to organs and tissues. The movement is directed by the blood flow which is calculated in each node using the Hagen-Poiseuille equation and Kirchhoff’s laws assuming laminar flow. Organs are voxelized and movement of cells from artery entry to vein exit is generated via a biased 3D random walk. The probabilities of whether cells move or stay in tissues are derived from rate constants of physiologically based compartment modeling. PET listmode data is generated using the Monte-Carlo simulation framework GATE based on the definition of a large-body PET scanner with cell paths as moving radioactive sources and the XCAT phantom providing attenuation data. From the flow simulation of 100,000 cells, 100 sample cells were further processed by GATE and listmode data was reconstructed into images for comparison. As demonstrated by comparisons of simulated and reconstructed cell distributions, CeFloPS has the capacity to simulate the cell behavior of whole-body PET. It achieves this simulation in a way that is anatomically and physiologically reasonable, thereby providing valuable data for the development and validation of cell tracking algorithms.
[…]
IEEEexample:BSTcontrol
@@ 1 Introduction
Cell imaging and tracking is a subject of active research gathering great interest in the medical and biomedical science community. Non-invasive imaging techniques using fluorescent or luminescent markers in optical imaging (OI), iron oxide particles in magnetic resonance imaging (MRI), or radio-labeled tracers in single photon emission computed tomography (SPECT) and positron emission tomography (PET) have been frequently used in preclinical studies to permit in vivo visualization of living cells [1]. For human application, MRI, PET and SPECT are the most promising tools as they provide either high spatial resolution (MRI) or excellent whole-body sensitivity (PET, SPECT). Of these imaging modalities, PET has become of special interest following labeled immune cells in the context of stem cell imaging and cell-based therapies [2].
@@
PET is particularly advantageous because it combines exceptional sensitivity with the possibility of whole-body coverage which has been further advanced in recent years by the design of large-field-of-view (LAFOV) PET systems with extended coverage of the axial field-of-view in the range of 1 to 2 meters, improved time-of-flight capabilities and excellent system sensitivities [3]. Using either long-lived isotopes or reporter genes, immunoPET has great potential to identify biomarkers allowing to treat patients based on their most likely response to a dedicated immunotherapy [4, 5, 6].
The idea to image or track biological cells with PET is highly related to the field of Positron Emission Particle Tracking (PEPT) which has been developed primarily for technical applications, such as evaluation of particle–fluid or particle–particle interactions in the pharmaceutical, chemical, oil, mining, construction and power generation industries [7]. The idea is to follow one or a few radio-labeled particles by tracking the coincidence lines-of-response (LORs) recognized from its two back-to-back gamma rays emitted after beta-plus decay. Thereby, a trajectory can be calculated by connecting LORs in a most probable way. As PEPT applications are generally carried out outside a clinical environment, special industrially usable devices were developed in parallel to the invention of clinical PET systems [8, 9]. Following the idea of PEPT in a clinical or pre-clinical environment, the outstanding sensitivity of PET can be utilized to track radio-labeled singular cells traveling inside a biological system (e.g. human, mouse) [9, 10, 11, 12].
The here presented work describes two simulations that build a foundation for development and validation of cell tracking in humans. The first simulation mimics the movement of radio-labeled biological cells in blood and their accumulation in tissues of the human body. To this end, the XCAT phantom was used to define the hemodynamic parameters for vascular blood flow, accompanied by a random walk approach to mimic the cell motion in the vascular bed [13]. Transition probabilities for the accumulation of cells in tissue were implemented via a tracer kinetic model approach. An XCAT-based approach has been used before to simulate blood flow in the vascular tree [14, 15]. The simulation of the transition of cells into the tissue and back via tracer kinetic models represents a new extension by us. Tracer kinetic models are commonly utilized in dynamic PET studies to measure how substances are distributed, metabolized, and eliminated within the body [16, 17]. Vice versa, these model parameters can also be utilized to simulate cell transitions.
The second simulation enables to generate raw PET data from the simulated cell pathways using Monte Carlo based physics simulations. Such simulations were also used in the PEPT context based on the Geant4 Application for Tomographic Emission (Gate) [18]. E.g. Herald et al. simulated both the ADAC/Phillips Forte and the Siemens Inveon pre-clinical PET and compared fluid simulations with experimental data [19, 20]. The foundation of our Gate simulation, developed here, is the concept of new total-body PET with an extensive field of view and high sensitivity, which presents novel possibilities for human whole-body cell tracking.
The task of cell tracking is not part of this work, but has been developed in many variants by different groups in the past [21, 5]. Nguyen et al. recently presented a new algorithm, PEPT-EM, which was able to successfully track the fate of more than 70 radio-labeled melanoma cells in a mouse model [12]. In this context, the group recently introduced an image reconstruction algorithm for PET that employs an optimal transport-based regularisation technique. This algorithm demonstrates considerable potential for the tracking of radioactively labelled cells within a human body, with no limitation on the number of trackable cells, and with the capacity to address the inherent handling of scattered, random and background events. The proposed methodology integrates the information from all detected events to facilitate the reconstruction of dynamic evolution, whilst concurrently enhancing the reconstruction quality through the enforcement of temporal consistency. [22, 23, 24].
@@
This study provides a unique workflow for cell flow PET simulation (CeFloPS) to simulate the movement of radio-labeled cells within the human body, serving as ground-truth data for developing and validating novel image reconstruction and cell tracking tools.
@@ 2.1 Simulation basis
CeFloPS is based on anatomical data provided by the XCAT phantom with high resolution 3D models of all organs, blood vessels, bones and muscles (Fig. 1a) in the form of STEP files [25]. Using Inventor (Version 2023, Autodesk Inc., San Rafael, USA), these STEP files were converted into StL files (StereoLithography file format), providing triangulated surface meshes of all organs and structures. The StL files were further processed in Python (Version 3.11.5, Python Software Foundation) with the help of the trimesh library [26], allowing for operations on the meshes and giving access to the coordinates of the vertices. For further processing, meshes of all muscles and internal organs were converted into voxelized volumes with voxel size 0.5 mm $\times$ 0.5 mm $\times$ 0.5 mm, while the complex anatomy of blood vessel meshes was simplified by reducing it to node networks. The nodes describe the possible paths a cell can take and are placed along the central axis of a blood vessel, storing information about (i) its spatial coordinates, (ii) which nodes are adjacent in the direction of blood flow and (iii) the radius of the blood vessel.
@@
Figure 1: Mesh representation of the XCAT phantom defining vessel structures and organs (a). Only the larger vessels are defined by XCAT showing partly missing connections, exemplarily shown on the arteries in the brain (b) and a just coarsely defined capillary bed, here shown for the kidney (c).
@@ 2.2 Blood flow simulation
$p_{\mathrm{node},i}=\frac{\overline{Q}_{i}}{\overline{Q}_{\mathrm{link}}}\quad\mathrm{with}\quad\overline{Q}_{\mathrm{link}}=\sum_{i=1}^{n}\overline{Q}_{i}$ (5)
@@
$\biggl{|}\frac{\mathrm{d}Q(x)}{\mathrm{d}x}\biggr{|}=\pi r^{2}\biggl{|}\frac{\mathrm{d}v(x)}{\mathrm{d}x}\biggr{|}=2Q_{\mathrm{max}}\frac{x}{r^{2}}\,,$ (7)
with $Q_{\mathrm{max}}=2\overline{Q}$. Division by the maximum blood flow rate yields the probability density function (PDF)
$f_{p}(x)=\frac{1}{Q_{\mathrm{max}}}\biggl{|}\frac{\mathrm{d}Q(x)}{\mathrm{d}x}\biggr{|}=2\frac{x}{r^{2}}\ \mathrm{with}\ \int_{0}^{r}f_{p}(x)\mathrm{d}x=1\,,$ (8)
@@
$F_{p}(x)=\int_{0}^{x}f_{p}(t)\mathrm{d}t=\frac{x^{2}}{r^{2}}\,.$ (9)
@@ 2.3 Organ distribution simulation
The random walk is initiated at a start-voxel, where a cell can move to one of the surrounding voxels. This gives 26 possible directions (provided the start-voxel does not reside on the organ’s periphery), all with the same probability. For the proceeding steps, the direction is biased in accordance with the previously determined weightings of the direction vectors in each voxel. If a cell reaches an end-voxel, the random walk concludes, followed by the cell exiting via the vein back into the nodal network for further blood flow simulation. The speed at which cells perform the random walk in the different organs is calibrated using the delay times determined by [30]. For this purpose, first the distribution of required time to travel from start- to end-voxel in each organ is determined for cells at an arbitrary random walk speed. The mean value of this distribution is calculated and compared with the value derived by [30], leading to new speed values for the corresponding organ.
@@
In addition to the biased random walk, a compartment model was incorporated into the organ distribution simulation to account for the transition state of a cell in tissue. Besides being spatially located in a voxel, a cell is also in a compartment state, which dictates whether a cell can move or not. The exchange between compartments is determined by transition rates $k_{i}$. The number of compartments and the values of the transition rates must be known in advance to the CeFloPS simulation and can be obtained from (physiologically based) kinetic models. As a first example and as proof-of-principle, we employ a two-compartment model comprising the compartments $C_{1}$ and $C_{2}$ (shown in Fig. 4). The compartment $C_{1}$ signifies a reversible binding of the cells, while $C_{2}$ denotes an irreversible binding. Consequently, if a cell is situated within either $C_{1}$ or $C_{2}$, it does not execute a random walk and remains temporarily or permanently in the same voxel. The change in amount of cells in these compartments can be described with the transition rates $K_{1}$, $k_{2}$, and $k_{3}$ through the differential equations
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}C_{1}(t)&=K_{1}C_{A}(t)-(k_{2}+k_{3})C_{1}(t)\quad\mathrm{and}\\ \frac{\mathrm{d}}{\mathrm{d}t}C_{2}(t)&=k_{3}C_{1}(t)\,.\end{split}$ (10)
The arterial input function (AIF), denoted by $C_{A}(t)$, serves as a metric for the concentration or number of cells present in the arterial blood and also has to be entered as input into the CeFloPS simulation. One method for generating an AIF is outlined in [31]. It is important to note that, in general, transition rates are only defined in the underlying kinetic model for a subset of the tissues included in CeFloPS. Consequently, the sum of all cells transitioning from the arterial blood into the specified tissues may not be equal to the rate of change of the AIF. In CeFlops, this would result in an erroneous distribution of cells, which would not correspond to (10). To counteract this, two virtual compartments were introduced. The first compartment, denoted by $C_{\mathrm{inj}}$, signifies the injection of cells into the arterial blood during the beginning of the simulation. The second compartment, designated by $C_{\mathrm{rem}}$, describes cells which are removed from the arterial blood and go into a tissue not defined in the kinetic model rates. The rate of change of the AIF is then given by
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}C_{A}(t)&=k_{\mathrm{inj}}(t)C_{\mathrm{inj}}(t)-k_{\mathrm{rem}}(t)C_{A}(t)\\ &+\sum_{\mathrm{tissues}}\left[k_{2}C_{1}(t)-K_{1}C_{A}(t)\right]\,.\end{split}$ (11)
The transition rates $k_{\mathrm{inj}}$ and $k_{\mathrm{rem}}$ are selected as time-dependent to account for the time-limited injection phase of the cells at the beginning of a simulation and potential delays during the transition into tissues. The determination of $k_{\mathrm{inj}}(t)$ and $k_{\mathrm{rem}}(t)$ is achieved through the implementation of a 3-layer neural network. The neural network identifies the optimal values for $k_{\mathrm{inj}}$ and $k_{\mathrm{rem}}$ at each simulation time $t$ by fitting the cell distributions in all tissues to fulfill the differential equations from (10) and (11). The latter phase of the distributions, in which a state of equilibrium is ultimately achieved, is given greater weight relative to the initial phase, during which significant alterations may occur. This ensures that the equilibrium phase at late stages of the simulation is mainly determined by the compartment model. The neural network is set up using the Python library PyTorch [32].
@@
Figure 4: The movement of cells within each tissue is regulated by compartmental models, which determine the probability of a cell undergoing a random walk or being constrained to a compartment.
The transition rates $k_{i}$ describe the transition behavior exhibited by the entirety of the cell population. However, in the CeFloPS simulation, each cell is considered independently. To determine the transition probability between compartments in a tissue for a single cell, the compartments are treated as states of a continuous Markov chain with a $3\!\times\!3$ transition rate matrix $\mathbf{G}$. The elements $\gamma_{i,j}$ of that matrix denote the transition rate from state $i$ to state $j$. Diagonal elements of $\mathbf{G}$ are determined by the condition that the row-sums of the matrix are equal to 0, i.e. $\sum_{j=1}^{3}\gamma_{i,j}=0$ for $i=[1,2,3]$. For the two-compartment model with rate constants $K_{1}$, $k_{2}$ and $k_{3}$, this results in the following matrix:
$\mathbf{G}=\begin{pmatrix}-K_{1}&K_{1}&0\\ k_{2}&-k_{2}-k_{3}&k_{3}\\ 0&0&0\end{pmatrix}\,.$
As shown in [33, 34, 35], one can derive a matrix $\mathbf{P}(t)$ of transition probabilities from $\mathbf{G}$, with elements $\pi_{i,j}(t)$ giving the transition probability from state $i$ to state $j$ during the length of a time step $t$:
$\mathbf{P}(t)=\exp(t\mathbf{G})=\sum_{n=0}^{\infty}\frac{t^{n}}{n!}\mathbf{G}^{n}\,,$ (12)
where $\exp$ is the matrix exponential. It takes into account that more than one transition can occur during a time step. Therefore the probabilities for transitions from $C_{A}$ to $C_{2}$ (i.e. $\pi_{1,3}$) are greater than 0, although in the two-compartment model no direct transition can occur between those compartments. In our simulations (12) is solved numerically using the Python package scipy [36].
In this example, we implement the rate constants of a 18F-Fluordesoxyglucose (FDG) model taken from [37] and listed in Table LABEL:tab:k_values. The application of kinetic modeling with FDG has been comprehensively addressed in the extant literature, thereby facilitating a more robust validation of our simulation compared to kinetic modeling with cells, which is still scarce in the literature and usually only takes few tissues into account. The simulated cells in this example therefore do not behave like real, existing cells, but like an FDG tracer. In this context, $C_{1}$ and $C_{2}$ denote the unbound and bound states of an FDG ”cell” in tissue and $C_{A}$ the free FDG ”cell” in the blood supply of the tissue. Based on the probabilities calculated with (12) from these transition rates, a random decision is made at each time step as to whether a cell transitions between these compartments. Fig. 5 demonstrates the temporal cell distributions calculated by the Markov chain approach in comparison to the whole population based compartment modelling using differential equations from (10), showing no systematic differences.
@@
Table 1: Values of some rate constants taken from [37]a and [38]b used in the simulation
@@
lunga 0.023 0.205 0.001
livera 0.660 0.765 0.002
spleena 1.593 2.867 0.006
pancreasb 0.3561 1.7077 0.0787
kidneysb 0.7023 1.3542 0.1778
Figure 5: Example model of cell distributions over time in blood, liver, and spleen after an injection at $t=0$ into the bloodstream. The cell dynamics were modeled using a (kinetic) compartment model based on (10) (solid lines) and a Markov chain approach based on (12) (dashed lines).
For the FDG simulation, the paths and distributions of 100,000 cells are simulated in parallel. For every time step, the organ and compartment in which each cell is located is recorded. This allows the generation of time profiles of cells in tissues. For a given tissue, the tissue concentration of cells per milliliter or the corresponding cell number (as the sum in $C_{1}$ and $C_{2}$) is calculated. The number of cells in the blood is made up of all cells that are still in the node network or in a $C_{A}$ compartment of an organ.